# Decentralized Stochastic Subgradient-based Methods for Nonsmooth Nonconvex Optimization

### v2

@@ 3.1 Basic assumptions
    There exists a locally Lipschitz continuous Lyapunov function $\psi:\mathbb{R}^{m}\to\mathbb{R}$ for the differential inclusion (3.2) with stable set $\mathcal{A}$.
@@
    The set $\{\psi({\bm{z}}):{\bm{z}}\in\mathcal{A}\}$ has empty interior in $\mathbb{R}$.
Assumption 3.3(1) clarifies the Lyapunov properties of a differential inclusion with respect to $-\Phi$, which is standard in the literature and can be satisfied by most appropriately selected set-valued mappings. Assumption 3.3(2) is the Weak Sard’s theorem, which stipulates that the set of values of $\psi$ at points outside the stable set is dense in $\mathbb{R}$. As demonstrated in demonstrated in [29, 37], Assumption 3.3(2) is mild in practice.
3.2 Global convergence
In this subsection, we establish the convergence properties of our proposed framework (DSM), under Assumption 3.1 and Assumption 3.3.
Towards this end, we first conduct consensus analysis of $\{{\bm{Z}}_{k}\}$. Denote ${\bm{P}}=\frac{1}{d}{\bm{1}_{d}}{\bm{1}}_{d}^{\top}$ and ${\bm{P}}_{\perp}={\bm{I}}_{d}-\frac{1}{d}{\bm{1}_{d}}{\bm{1}}_{d}^{\top}$, then any matrix ${\bm{Z}}_{k}\in\mathbb{R}^{m\times d}$ has the orthogonal decomposition ${\bm{Z}}_{k}={\bm{Z}}_{k}{\bm{P}}+{\bm{Z}}_{k}{\bm{P}}_{\perp}$. Define the disagreement matrix ${\bm{Z}}_{\perp,k}:={\bm{Z}}_{k}{\bm{P}}_{\perp}$. Then the above orthogonal decomposition can be rephrased as ${\bm{Z}}_{k}={\bm{Z}}_{k}{\bm{P}}+{\bm{Z}}_{\perp,k}$.
We begin our proof with the following preliminary lemma from [53, Lemma 3.1].
Lemma 3.4.
Let $\{\gamma_{k}\}$ be a sequence of real numbers. If $10<\alpha<$ 1 and $\lim_{k\rightarrow+\infty}\gamma_{k}=\gamma$, then it holds that $\lim_{k\rightarrow+\infty}\sum_{\ell=0}^{k}\alpha^{k-\ell}\gamma_{\ell}=\frac{\gamma}{1-\alpha}$.
Next, we present the following lemma to show that $\left\|\frac{1}{d}({\bm{Z}}_{k+1}-{\bm{Z}}_{k}){\bm{1}}_{d}\right\|$ converges to 00 as $k$ goes to infinity.
Lemma 3.5.
Suppose Assumption 3.1 holds, then for any $\{{\bm{Z}}_{k}\}$ generated by the framework (DSM), we have $\lim_{k\to+\infty}\left\|\eta_{k}({\bm{H}}_{k}+\Xi_{k+1})\right\|=0$. Moreover, it holds that $\lim_{k\to+\infty}\left\|\frac{1}{d}({\bm{Z}}_{k+1}-{\bm{Z}}_{k}){\bm{1}_{d}}\right\|=0$.
@@ 3.2 Global convergence
From Assumption 3.1(1) and the fact that $\lim_{k\to+\infty}\eta_{k}=0$ in Assumption 3.1(2), it holds that $\limsup_{k\to+\infty}\eta_{k}\left\|{\bm{H}}_{k}\right\|=0$. Furthermore, Assumption 3.1(2) illustrates that
$\limsup_{i\to+\infty}\left\|\eta_{i}\Xi_{i+1}\right\|\leq\lim_{i\to+\infty}\sup_{i\leq j\leq\Lambda_{\eta}(\lambda_{\eta}(i)+T)}\left\|\sum_{k=i}^{j}\eta_{k}\Xi_{k+1}\right\|=0.$
Therefore, we can conclude that
$\lim_{k\to+\infty}\left\|\eta_{k}({\bm{H}}_{k}+\Xi_{k+1})\right\|=0.$ (3.3)
Moreover, from the update scheme of $\{{\bm{Z}}_{k}\}$ in (DSM) and Definition 2.1, it holds that $\frac{1}{d}{\bm{Z}}_{k+1}{\bm{1}_{d}}=\frac{1}{d}{\bm{Z}}_{k}{\bm{1}_{d}}-\frac{\eta_{k}}{d}({\bm{H}}_{k}+\Xi_{k+1}){\bm{1}_{d}}$. Together with (3.3), we have
$\lim_{k\to+\infty}\frac{1}{d}\left\|({\bm{Z}}_{k+1}-{\bm{Z}}_{k}){\bm{1}}_{d}\right\|=\lim_{k\to+\infty}\left\|\frac{\eta_{k}}{d}({\bm{H}}_{k}+\Xi_{k+1}){\bm{1}}_{d}\right\|$
[…]
This completes the proof. ∎
Intuitively, ${\bm{Z}}_{\perp,k}$ measures the dissimilarity of local variables of all the agents at $k$-th iteration. In the following proposition, we prove that $\lim_{k\to+\infty}{\bm{Z}}_{\perp,k}=0$, i.e., the dissimilarity of local variables of all the agents converges to 00.
Proposition 3.6.
Suppose Assumption 3.1 holds, then for any $\{{\bm{Z}}_{k}\}$ generated by the framework (DSM), it holds that
$\lim_{k\to+\infty}\left\|{\bm{Z}}_{\perp,k}\right\|=0.$ (3.4)
Proof.
From Definition 2.1, we can conclude that there exists $\alpha\in(0,1]$ such that $\left\|{\bm{W}}({\bm{I}}_{d}-{\bm{P}})\right\|=1-\alpha$. Then for any $k\geq 0$, it follows from the update scheme of $\{{\bm{Z}}_{k}\}$ in (DSM) that
$\left\|{\bm{Z}}_{\perp,k+1}\right\|=\left\|[{\bm{Z}}_{k}{\bm{W}}-\eta_{k}({\bm{H}}_{k}+\Xi_{k+1})]({\bm{I}}_{d}-{\bm{P}})\right\|$
@@
Here the second inequality follows from the fact that $({\bm{I}}_{d}-{\bm{P}}){\bm{W}}({\bm{I}}_{d}-{\bm{P}})={\bm{W}}({\bm{I}}_{d}-{\bm{P}})$. Since Lemma 3.5 illustrates that $\lim_{k\to+\infty}\left\|\eta_{k}({\bm{H}}_{k}+\Xi_{k+1})\right\|=0$, then it can be readily deduced by Lemma 3.4 that $\lim_{k\to+\infty}\left\|{\bm{Z}}_{\perp,k}\right\|=0$. This completes the proof. ∎
As illustrated in Proposition 3.6, any sequence $\{{\bm{Z}}_{k}\}$ from (DSM) satisfies $\lim_{k\to+\infty}\left\|{\bm{Z}}_{k}-{\bm{Z}}_{k}{\bm{P}}\right\|=0$. As a result, we can conclude that the cluster points of the sequence $\{{\bm{Z}}_{k}\}$ are exactly the same as those of the sequence $\{{\bm{Z}}_{k}{\bm{P}}\}$.
To analyze the convergence properties of $\{\frac{1}{d}{\bm{Z}}_{k}{\bm{1}}_{d}\}$, we present the following proposition to illustrate the update scheme of $\{\frac{1}{d}{\bm{Z}}_{k}{\bm{1}}_{d}\}$.
Proposition 3.7.
Suppose Assumption 3.1 holds, then there exists a nonnegative diminishing sequence $\{\tilde{\delta}_{k}\}$ such that for any sequence $\{{\bm{Z}}_{k}\}$ generated from the framework (DSM), it holds for all $k\geq 0$ that
$\frac{1}{d}{\bm{Z}}_{k+1}{\bm{1}}_{d}\in\frac{1}{d}{\bm{Z}}_{k}{\bm{1}}_{d}-\eta_{k}\left(\Phi^{\tilde{\delta}_{k}}(\frac{1}{d}{\bm{Z}}_{k}{\bm{1}}_{d})+\frac{1}{d}\Xi_{k+1}{\bm{1}}_{d}\right).$ (3.5)
@@
From the update scheme of $\{{\bm{Z}}_{k}\}$ in (DSM), we can conclude that
$\frac{1}{d}{\bm{Z}}_{k+1}{\bm{1}}_{d}$ $=\frac{1}{d}{\bm{Z}}_{k}{\bm{W}}{\bm{1}}_{d}-\frac{\eta_{k}}{d}({\bm{H}}_{k}+\Xi_{k+1}){\bm{1}}_{d}$
$=\frac{1}{d}{\bm{Z}}_{k}{\bm{1}}_{d}-\frac{\eta_{k}}{d}{\bm{H}}_{k}{\bm{1}}_{d}-\frac{\eta_{k}}{d}\Xi_{k+1}{\bm{1}}_{d}.$
Let $\delta^{\star}_{k}=\left\|{\bm{Z}}_{k}{\bm{P}}_{\perp}\right\|$, from the definition of ${\bm{P}}_{\perp}$, it holds that $\left\|{\bm{z}}_{i,k}-\frac{1}{d}{\bm{Z}}_{k}{\bm{1}}_{d}\right\|\leq\left\|{\bm{Z}}_{k}{\bm{P}}_{\perp}\right\|=\delta^{\star}_{k}$.
For simplicity, we denote $\mathrm{conv}\,(\frac{1}{d}\sum_{i=1}^{d}\Phi_{i}^{\delta_{k}+\delta^{\star}_{k}}(\frac{1}{d}{\bm{Z}}_{k}{\bm{1}}_{d}))$ by $\mathcal{C}_{k}$. Together with Assumption 3.1(3), we can conclude that
$\frac{1}{d}{\bm{H}}_{k}{\bm{1}}_{d}\in\mathrm{conv}\,\left(\frac{1}{d}\sum_{i=1}^{d}\Phi_{i}^{\delta_{k}}({\bm{z}}_{i,k})\right)\subseteq\mathcal{C}_{k}.$ (3.6)
Now we claim that there exists a nonnegative diminishing sequence $\{\tilde{\delta}_{k}\}$ such that
$\mathcal{C}_{k}\subseteq\Phi^{\tilde{\delta}_{k}}(\frac{1}{d}{\bm{Z}}_{k}{\bm{1}}_{d}).$ (3.7)
Otherwise, there exists $\delta_{\varepsilon}>0$ and a subsequence $\{k_{j}\}\subset\mathbb{N}_{+}$ such that
$\sup\left\{\mathrm{dist}\left({\bm{y}},\Phi^{\delta_{\varepsilon}}(\frac{1}{d}{\bm{Z}}_{k_{j}}{\bm{1}}_{d})\right):{\bm{y}}\in\mathcal{C}_{k_{j}}\right\}>0.$ (3.8)
From the uniform boundedness of $\{{\bm{Z}}_{k}\}$ in Assumption 3.1(1), without loss of generality, we can assume that $\{{\bm{Z}}_{k_{j}}\}$ converges to $\tilde{{\bm{Z}}}\in\mathbb{R}^{n\times d}$. Then from the graph closedness and local boundedness of $\Phi_{i}$, we have
$\lim_{j\to+\infty}\sup\left\{\mathrm{dist}\left({\bm{y}},\Phi_{i}(\frac{\tilde{{\bm{Z}}}{\bm{1}}_{d}}{d})\right):{\bm{y}}\in\Phi_{i}^{\delta_{k_{j}}+\delta^{\star}_{k_{j}}}(\frac{{\bm{Z}}_{k_{j}}{\bm{1}}_{d}}{d})\right\}=0.$
Then based on Jensen’s inequality, it holds that
$\lim_{j\to+\infty}\sup\left\{\mathrm{dist}\left({\bm{y}},\Phi(\frac{1}{d}\tilde{{\bm{Z}}}{\bm{1}}_{d})\right):{\bm{y}}\in\mathcal{C}_{k_{j}}\right\}=0.$
which contradicts (3.8). As a result, there exists a nonnegative diminishing sequence $\{\tilde{\delta}_{k}\}$ such that (3.7) holds, hence
$\frac{1}{d}{\bm{Z}}_{k+1}{\bm{1}}_{d}$ $=\frac{1}{d}{\bm{Z}}_{k}{\bm{W}}{\bm{1}}_{d}-\eta_{k}({\bm{H}}_{k}+\frac{1}{d}\Xi_{k+1}){\bm{1}}_{d}$ (3.9)
$\in\frac{1}{d}{\bm{Z}}_{k}{\bm{1}}_{d}-\eta_{k}\left(\Phi^{\tilde{\delta}_{k}}(\frac{1}{d}{\bm{Z}}_{k}{\bm{1}}_{d})+\frac{1}{d}\Xi_{k+1}{\bm{1}}_{d}\right).$
This completes the proof. ∎
Proposition 3.8.
Suppose Assumption 3.1 holds. Then the interpolated process of the sequence $\{\frac{1}{d}{\bm{Z}}_{k}{\bm{1}}_{d}\}$ is a perturbed solution to the differential inclusion (3.2).

### v3

@@ 3.1 Basic assumptions and main results
    The sequence of step-sizes $\{\eta_{k}\}$ is positive and satisfies $\sum_{k=0}^{\infty}\eta_{k}=+\infty$.
@@
    The sequences $\{{\bm{H}}_{k}\}$ and $\{\Xi_{k+1}\}$ are uniformly bounded in $\mathbb{R}^{m\times d}$, whenever $\{{\bm{Z}}_{k}\}$ is bounded in $\mathbb{R}^{m\times d}$.
  3. 3.
    There exists a locally Lipschitz continuous and coercive Lyapunov function $\psi:\mathbb{R}^{m}\to\mathbb{R}$ for the differential inclusion (3.1) with stable set $\mathcal{A}$.
  4. 4.
    The set $\{\psi({\bm{z}}):{\bm{z}}\in\mathcal{A}\}$ is a finite subset of $\mathbb{R}$.
Assumption 3.1(1) allows for a flexible choice of the step-size, while Assumption 3.1(2) is reasonable and mild, which is commonly employed in various existing works [36, 37, 43, 40, 38].
Assumption 3.1(3) clarifies the Lyapunov properties of a differential inclusion with respect to $-\Phi$, which is standard in the literature [45, 65, 66, 37, 36, 63] and can be satisfied by most appropriately selected set-valued mappings. Since each $\Phi_{i}$ is graph-closed and locally bounded, the convex-valued set-valued mapping $\Phi$ is also graph-closed and locally bounded. Then, the coercivity of $\psi$ in Assumption 3.1(3) indicates the compactness of any level set $\mathcal{L}_{r}:=\{x\in\mathbb{R}^{n}:\psi(x)\leq r\}$.
Assumption 3.1(4) is similar to the Weak Sard’s condition in [37], which stipulates that the set of values of $\psi$ at points outside the stable set is dense in $\mathbb{R}$. As demonstrated in [36], Assumption 3.1(4) holds when $f$ is a definable function with the selection $\Phi:=\partial f$ and $\psi:=f$.
In the following, we present some basic notations and definitions throughout Section 3. For any $M>0$, the stopping time $\tau_{M}$ is defined as
$\tau_{M}:=\inf\{k\in\mathbb{N}:\|{\bm{Z}}_{k}\|>M\},$
and the upper bound of noisy update direction before stopping time is given by
$\ell_{M}:=\sup_{0\leq k\leq\tau_{M}}\left\|{\bm{H}}_{k}+\Xi_{k+1}\right\|,$
which is finite followed by Assumption 3.1(2). The second largest singular value of mixing matrix ${\bm{W}}$ is
$\lambda_{2}:=\left\|{\bm{W}}({\bm{I}}_{d}-\frac{{\bm{1}}_{d}{\bm{1}}_{d}^{\top}}{d})\right\|.$
Definition 3.2.
For any positive sequence $\{\eta_{k}\}$, we say $\{\eta_{k}\}$ is $\alpha_{\text{ub}}$-upper-bounded or is upper-bounded by $\alpha_{\text{ub}}$, if $\sup_{k\in\mathbb{N}}\eta_{k}\leq\alpha.$
Definition 3.3.
For any sequence of vectors (or matrices) $\{\xi_{k}\}$, and given constants $T,\alpha_{\text{ub}}>0$, we say $\{\xi_{k}\}$ is $(\alpha_{\text{ub}},T,\{\eta_{k}\})$-controlled, if for any $s\geq 0$, we have
$\sup_{s\leq i\leq\Lambda_{\eta}(\lambda_{\eta}(s)+T)}\left\|\sum_{k=s}^{i}\eta_{k}\xi_{k+1}\right\|\leq\alpha_{\text{ub}}.$
Moreover, we say $\{\xi_{k}\}$ is $(\alpha_{\text{ub}},\alpha,T,\{\eta_{k}\})$-asymptotically controlled, if it is $(\alpha_{\text{ub}},T,\{\eta_{k}\})$-controlled and
$\limsup_{s\to\infty}\sup_{s\leq i\leq\Lambda_{\eta}(\lambda_{\eta}(s)+T)}\left\|\sum_{k=s}^{i}\eta_{k}\xi_{k+1}\right\|\leq\alpha.$
With above tools in placement, we first present Proposition 3.4 to demonstrate the average iterative sequence of (DSM) before stopping time coincides with the update scheme (2.3).
Proposition 3.4.
Suppose Assumption 3.1 holds, and let $\mathcal{Z}_{0}$ be a compact subset of $\mathbb{R}^{m}$. The sequence $\{{\bm{Z}}_{k}\}$ is generated by (DSM), and ${\bm{Z}}_{0}={\bm{z}}_{0}{\bm{1}}_{d}^{\top},{\bm{z}}_{0}\in\mathcal{Z}_{0}$. Then the recursion relation of $\{\frac{1}{d}{\bm{Z}}_{k\land\tau_{M}}{\bm{1}}_{d}\}$ can be reformulated as
$\frac{1}{d}{\bm{Z}}_{(k+1)\land\tau_{M}}{\bm{1}}_{d}\in\frac{1}{d}{\bm{Z}}_{k\land\tau_{M}}{\bm{1}}_{d}-\eta_{k}[\Phi^{s_{k}}(\frac{1}{d}{\bm{Z}}_{k\land\tau_{M}}{\bm{1}}_{d})+\frac{1}{d}\Xi_{k+1}{\bm{1}}_{d}]\mathbb{1}_{\tau_{M}>k},$ (3.2)
where $s_{k}:=\delta_{k}+\|{\bm{Z}}_{\perp,k\land\tau_{M}}\|\leq\delta_{k}+\frac{\sup_{0\leq i\leq k\land\tau_{M}}\eta_{i}\ell_{M}}{1-\lambda_{2}}$.
@@
From (DSM), the update scheme for $\{{\bm{Z}}_{k\land\tau_{M}}\}$ can be expressed as
${\bm{Z}}_{(k+1)\land\tau_{M}}={\bm{Z}}_{k\land\tau_{M}}+[{\bm{Z}}_{k\land\tau_{M}}({\bm{W}}-{\bm{I}}_{d})-\eta_{k}({\bm{H}}_{k}+{\Xi}_{k+1})]{\mathbb{1}_{\tau_{M}>k}},$ (3.3)
where $\mathbb{1}_{\tau_{M}>k}$ is the indicator function for the event $\{\tau_{M}>k\}$. Alternatively, the update can be written as
${\bm{Z}}_{(k+1)\land\tau_{M}}={\bm{Z}}_{k\land(\tau_{M}-1)}{\bm{W}}-\eta_{k\land(\tau_{M}-1)}({\bm{H}}_{k\land(\tau_{M}-1)}+{\Xi}_{(k+1)\land\tau_{M}}).$ (3.4)
Define the consensus projection matrix ${\bm{P}}:=\frac{1}{d}{\bm{1}_{d}}{\bm{1}}_{d}^{\top}$ and the disagreement matrix ${\bm{Z}}_{\perp,k}:={\bm{Z}}_{k}({\bm{I}}_{d}-{\bm{P}})$. Then, the orthogonal decomposition holds:
${\bm{Z}}_{k}={\bm{Z}}_{k}{\bm{P}}+{\bm{Z}}_{\perp,k}.$
Intuitively, ${\bm{Z}}_{\perp,k}$ measures the dissimilarity among all agents’ local variables at $k$-th iteration.
According to (3.4), simple calculations yield that
$\|{\bm{Z}}_{\perp,(k+1)\land\tau_{M}}\|={}$ $\left\|({\bm{Z}}_{k\land(\tau_{M}-1)}{\bm{W}}-\eta_{k\land(\tau_{M}-1)}({\bm{H}}_{k\land(\tau_{M}-1)}+{\Xi}_{(k+1)\land\tau_{M}}))({\bm{I}}_{d}-{\bm{P}})\right\|$ (3.5)
@@
[…]
where the second inequality follows from the fact that $({\bm{I}}_{d}-{\bm{P}}){\bm{W}}({\bm{I}}_{d}-{\bm{P}})={\bm{W}}({\bm{I}}_{d}-{\bm{P}})$ and the last inequality is due to ${\bm{Z}}_{\perp,0}={\bm{Z}}_{0}-\frac{{\bm{Z}}_{0}{\bm{1}}_{d}{\bm{1}}_{d}^{\top}}{d}=0$. As a result, we obtain that
$\|{\bm{Z}}_{\perp,l\land\tau_{M}}\|\leq\frac{\sup_{0\leq i\leq k\land\tau_{M}}\eta_{i}\ell_{M}}{1-\lambda_{2}},\text{ for any }l\leq k+1.$ (3.6)
We right-multiply $\frac{{\bm{1}}_{d}}{d}$ to (3.3) to get:
$\frac{1}{d}{\bm{Z}}_{(k+1)\land\tau_{M}}{\bm{1}}_{d}=\frac{1}{d}{\bm{Z}}_{k\land\tau_{M}}{\bm{1}}_{d}-\eta_{k}(\frac{1}{d}{\bm{H}}_{k}{\bm{1}}_{d}+\frac{1}{d}\Xi_{k+1}{\bm{1}}_{d})\mathbb{1}_{\tau_{M}>k}.$ (3.7)
Equation (1.5) reveals that $\frac{1}{d}{\bm{H}}_{k}{\bm{1}_{d}}\in\mathrm{conv}\,\left(\frac{1}{d}\sum_{i=1}^{d}\Phi_{i}^{\delta_{k}}({\bm{z}}_{i,k})\right)$. According to Carathéodory’s Theorem, it admits the representation
$\frac{1}{d}{\bm{H}}_{k}{\bm{1}_{d}}=\sum_{j=0}^{m}c_{j}(\frac{1}{d}\sum_{i=1}^{d}{\bm{y}}_{i,k}^{j}),\quad{\bm{y}}_{i,k}^{j}\in\Phi_{i}^{\delta_{k}}({\bm{z}}_{i,k}),c_{j}\in\Delta_{m}.$
On the event $\{\tau_{M}>k\}$, combining with (3.6) yields that
$~{}\mathrm{dist}\left(\frac{1}{d}{\bm{H}}_{k}{\bm{1}_{d}},\mathrm{conv}\,\left(\frac{1}{d}\sum_{i=1}^{d}\Phi_{i}(\frac{1}{d}{\bm{Z}}_{k\land\tau_{M}}{\bm{1}}_{d}+\mathbb{B}({\bm{0}},\delta_{k}+\|{\bm{Z}}_{\perp,k\land\tau_{M}}\|))\right)\right)$
$\leq$ $~{}\sum_{j=0}^{m}c_{j}{}\mathrm{dist}\left(\frac{1}{d}\sum_{i=1}^{d}{\bm{y}}_{i,k}^{j},\frac{1}{d}\sum_{i=1}^{d}\Phi_{i}\left(\frac{1}{d}{\bm{Z}}_{k\land\tau_{M}}{\bm{1}}_{d}+\mathbb{B}({\bm{0}},\delta_{k}+\|{\bm{Z}}_{\perp,k\land\tau_{M}}\|)\right)\right)$
$\leq$ $~{}\sum_{j=0}^{m}c_{j}\frac{1}{d}\sum_{i=1}^{d}\mathrm{dist}\left({\bm{y}}_{i,k}^{j},\Phi_{i}\left(\frac{1}{d}{\bm{Z}}_{k\land\tau_{M}}{\bm{1}}_{d}+\mathbb{B}({\bm{0}},\delta_{k}+\|{\bm{Z}}_{\perp,k\land\tau_{M}}\|)\right)\right)$
$\leq$ $~{}\sum_{j=0}^{m}c_{j}\frac{1}{d}\sum_{i=1}^{d}\delta_{k}=\delta_{k},$
where the first inequality is due to the Jensen’s inequality and the second inequality is followed by the triangle inequality. That is to say,
$\frac{1}{d}{\bm{H}}_{k}{\bm{1}_{d}}\in\Phi^{s_{k}}\left(\frac{1}{d}{\bm{Z}}_{k\land\tau_{M}}{\bm{1}}_{d}\right).$
Plugging this relation into (3.7) completes the proof. ∎
Based on Proposition 3.4, Theorem 3.5 establishes that the sequence $\{{\bm{Z}}_{k}\}$ remains uniformly bounded, whenever step-size $\{\eta_{k}\}$, $\{\delta_{k}\}$ are upper-bounded by a sufficiently small $\alpha_{ub}>0$, and $\{\Xi_{k+1}\}$ is $(\alpha_{ub},T,\{\eta_{k}\})$-controlled for some $T>0$.
Theorem 3.5.
Suppose Assumption 3.1 holds, and let $\mathcal{Z}_{0}$ be a compact subset of $\mathbb{R}^{m}$. Then for any $r>\max\{0,4\sup_{x\in\mathcal{Z}_{0}\cup\mathcal{A}}\psi(x)\}$, there exist $\alpha_{ub}>0,T>0$ such that for any $\alpha_{ub}$-upper-bounded sequences $\{\eta_{k}\}$ and $\{\delta_{k}\}$, and any $(\alpha_{ub},T,\{\eta_{k}\})$-controlled sequence $\{\Xi_{k+1}\}$, the sequence $\{{\bm{Z}}_{k}\}$ generated by (DSM) with ${\bm{Z}}_{0}={\bm{z}}_{0}{\bm{1}}_{d}^{\top},{\bm{z}}_{0}\in\mathcal{Z}_{0}$ is restricted in $\mathcal{L}_{2r}^{d}$.
@@
Recall from Proposition 3.4, we have
$\frac{1}{d}{\bm{Z}}_{(k+1)\land\tau_{M}}{\bm{1}}_{d}\in\frac{1}{d}{\bm{Z}}_{k\land\tau_{M}}{\bm{1}}_{d}-\eta_{k}[\Phi^{s_{k}}(\frac{1}{d}{\bm{Z}}_{k\land\tau_{M}}{\bm{1}}_{d})+\frac{1}{d}\Xi_{k+1}{\bm{1}}_{d}]\mathbb{1}_{\tau_{M}>k}.$
For any $M>0$, employing Lemma 2.15, we conclude that there exists $\alpha$, $T>0$ such that for any $\alpha$-upper-bounded sequences $\{\eta_{k}\}$ and $\{s_{k}\}$, and any $(\alpha,T,\{\eta_{k}\})$-controlled sequence $\{\frac{1}{d}\Xi_{k+1}{\bm{1}}_{d}\mathbb{1}_{\tau_{M}>k}\}$, the sequence $\{\frac{1}{d}{\bm{Z}}_{k\land\tau_{M}}{\bm{1}}_{d}\}$ is restricted in $\mathcal{L}_{r}$.
Let $L_{1}$ be the Lipschitz constant of $\psi$ in $\mathcal{L}_{r}+\mathbb{B}({\bm{0}},\frac{\alpha\ell_{M}}{1-\lambda_{2}})$, and $\alpha_{ub}:=\min\{\frac{\alpha}{2},\frac{\alpha(1-\lambda_{2})}{2\ell_{M}},\frac{r(1-\lambda_{2})}{L_{1}\ell_{M}}\}$. It can be readily verified that
$\psi({\bm{z}}_{i,k\land\tau_{M}})-\psi(\frac{1}{d}{\bm{Z}}_{k\land\tau_{M}}{\bm{1}}_{d})\leq L_{1}\frac{\sup_{0\leq i\leq k\land\tau_{M}}\eta_{i}\ell_{M}}{1-\lambda_{2}}\leq r,$
for any $i\in[d]$ and any $\alpha_{ub}$-upper-bounded $\{\eta_{k}\}$.
Therefore, for any $M>0$, any $\alpha_{ub}$-upper-bounded sequences $\{\eta_{k}\}$, $\{\delta_{k}\}$, and any $(\alpha_{ub},T,\{\eta_{k}\})$-controlled sequence $\{\Xi_{k+1}\}$, it holds that $\{\eta_{k}\}$ and $\{s_{k}\}$ are $\alpha$-upper-bounded, $\{\frac{1}{d}\Xi_{k+1}{\bm{1}}_{d}\mathbb{1}_{\tau_{M}>k}\}$ is $(\alpha,T,\{\eta_{k}\})$-controlled, thus $\{{\bm{z}}_{i,k\land\tau_{M}}\}$ is restricted in $\mathcal{L}_{2r},\forall i\in[d]$. Taking $M>\sup\{\|{\bm{Z}}\|:{\bm{Z}}\in\mathcal{L}^{d}_{2r}\}$, from the definition of the stopping time $\tau_{M}$, we can derive that $\tau_{M}=+\infty$, then ${\bm{z}}_{i,k}={\bm{z}}_{i,k\land\tau_{M}}\in\mathcal{L}_{2r},\forall i\in[d],k\geq 0$, which completes the proof. ∎
The following lemma from [53, Lemma 3.1] concerns the limit property of a convolution-like scalar sequence.
Lemma 3.6.
Let $\{\gamma_{k}\}$ be a sequence of real numbers. If $10<a<$ 1 and $\lim_{k\rightarrow\infty}\gamma_{k}=\gamma$, then it holds that $\lim_{k\rightarrow\infty}\sum_{\ell=0}^{k}a^{k-\ell}\gamma_{\ell}=\frac{\gamma}{1-a}$.
Combining Theorem 3.5 and Lemma 3.6, we demonstrate in Theorem 3.7 that the sequence $\{\bm{Z}_{k}\}$ asymptotically reaches consensus and converges to the stable set $\mathcal{A}$ with diminishing upper-bounded $\{\eta_{k}\}$, $\{\delta_{k}\}$ and $(\alpha_{ub},0,T,\{\eta_{k}\})$ asymptotically controlled noises.
Theorem 3.7.
Suppose Assumption 3.1 holds, and let $\mathcal{Z}_{0}$ be a compact subset of $\mathbb{R}^{m}$, $\{{\bm{Z}}_{k}\}$ be generated by (DSM) with ${\bm{Z}}_{0}={\bm{z}}_{0}{\bm{1}}_{d}^{\top},{\bm{z}}_{0}\in\mathcal{Z}_{0}$. Then there exist $\alpha_{ub}>0,T>0$ such that for any $\alpha_{ub}$-upper-bounded diminishing sequences $\{\eta_{k}\}$ and $\{\delta_{k}\}$, and any $(\alpha_{ub},0,T,\{\eta_{k}\})$-asymptotically controlled sequence $\{\Xi_{k+1}\}$, it follows that
$\lim_{k\to\infty}\mathrm{dist}({\bm{Z}}_{k},\{{\bm{Z}}\in\mathbb{R}^{m\times d}:{\bm{Z}}={\bm{z}}{\bm{1}}^{\top},{\bm{z}}\in\mathcal{A}\})=0,$
and $\psi({\bm{z}}_{i,k})$ converges for each $i\in[d].$